OpenStudy (anonymous):

Can someone check my answers? It's Advanced Algebra with Financial Applications. There are only 6 six questions. (:

OpenStudy (anonymous):

The number of customers that visit a local small business is 51,200 and has been continuously declining at a rate of 3.8% each year. What is the approximate number of customers that visit the business in 14 years? 33,690 29,766 <-- My answer 30,076 87,159

OpenStudy (anonymous):

Shelby purchased a plot of land for $27,500. The land appreciates about 2.9% each year. What is the value of the land after nine years? $21,101.13 $35,568.95 <-- My answer $34,677.50 $28,103.94

OpenStudy (anonymous):

Luke purchased a motorcycle for $8,765. It depreciates about 5.3% each year. What is the value of the motorcycle after five years? $6,442.27 $6,675.78 <-- My answer $8,573.14 $11,347.33

OpenStudy (anonymous):

The yearly attendance at a local movie theater is 82,000 and grows continuously at a rate of 5.1% each year. What is the approximate attendance at the movie theater in ten years? 136,554 85,552 134,847 <-- My answer 123,820

OpenStudy (thejax):

I think u got it right.

OpenStudy (anonymous):

A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be 3 interactions. If there are 3 children, there can be 12 interactions. If there are 4 children, there can be 39 interactions. Which recursive equation represents the pattern? an = an – 1 + 3(n – 1) <-- My answer (I completely guess on this one :l ) an = an – 1 + 3(n – 1) an = an – 1 + (3n – 1) an = an – 1 + (n – 1)3

OpenStudy (anonymous):

Last one (: A playground is being designed where children can interact with their friends in certain combinations. If there is 1 child, there can be 0 interactions. If there are 2 children, there can be 3 interactions. If there are 3 children, there can be 9 interactions. If there are 4 children, there can be 18 interactions. How many interactions will there be for 9 children? 135 84 <-- My answer (Again, this is just a guess :( ) 93 108

OpenStudy (thejax):

the progression is correct

OpenStudy (anonymous):

The progression? I'm sorry not sure which one that is

OpenStudy (thejax):

an = an – 1 + 3(n – 1)

OpenStudy (anonymous):

Okay, I figured that was the one you were talking about lol

OpenStudy (anonymous):

Do you think you could help me out with the last one then?

OpenStudy (anonymous):

Or can anyone try and help me out? I would really like to get an A on this :D

OpenStudy (thejax):

....

OpenStudy (anonymous):

Or not...? I'll just figure it out on my one then....

OpenStudy (anonymous):

own*

OpenStudy (tkhunny):

Business - Perfect Land - Perfect Motorcycle - Perfect Theatre - No Good. You missed the word "continuously". Not Continuous: \(82000\cdot 1.051^{10}\) = Your Answer Continuous: \(82000\cdot e^{0.051\cdot 10}\) = Correct Answer

OpenStudy (anonymous):

How would I work that out?

OpenStudy (anonymous):

Never mind, I figured it out. The yearly attendance at a local movie theater is 82,000 and grows continuously at a rate of 5.1% each year. What is the approximate attendance at the movie theater in ten years? 136,554 <-- My NEW answer 85,552 134,847 <-- My OLD answer 123,820

OpenStudy (anonymous):

@skullpatrol I can see you are the only one still looking at this, so can you help me out a little? If not, then please let me know (:

OpenStudy (anonymous):

Okay, well thanks to those that helped out and checked my work for me (: Closing the question now.